# Cryptographic Censorship

### 05 Feb 2024 v1

@@ 1 Introduction
Curvature singularities are ubiquitous in General Relativity: the evolution of smooth initial data generically results in such singularities [1, 2, 3]. Observationally, however, evidence for singularity formation from generic gravitational processes (such as matter collapse) remains at best indirect. This apparent absence of directly-observable singularities is often explained by the general expectation that the formation of horizons is likewise generic: so singularities are usually cloaked behind event horizons. This hypothesis — that singularities resulting from initial data evolution should be generically hidden behind horizons — is known as the Weak Cosmic Censorship Conjecture (WCCC) [4]. It is a cornerstone assumption of almost all results in modern gravitational physics, from classical theorems such as the Hawking area law [5], to key results from the general “it from qubit” program. In particular, the expectation that CFT thermalization corresponds to black hole equilibration, and more generally that black hole statistics are reflected in CFT thermal states, crucially depends on the assumption that states with horizons are typical in some measure. [^1] This is not identical to the firewall sense of typicality of [6]. See footnote 7 on page 7. If the WCCC were false, naked singularities could be as typical as black holes; our understanding of quantum chaos, eigenstate thermalization, and black hole information in holography would be called into question.
@@ A General Guarantee for Horizon Formation
More concretely, we ask when there is a guarantee for the existence of some operator $Q$ in the large-$N$ limit that cannot be reconstructed from knowledge of the causal wedge alone, given that the latter can be reconstructed efficiently. [^4] By an efficient reconstruction, we mean one that can be implemented by a quantum circuit whose size is polynomially bounded in the logarithm of the dimension of the code subspace under consideration.
In [32], two of us provided such a guarantee at large but finite $N$ under certain assumptions. The full statement of the relevant theorem will be reviewed at length in Sec. 3.1; here we provide a heuristic statement only, starting with the most significant assumption: that the action of the fundamental time evolution operator $U$ of the system is well modeled on our code subspace by a pseudorandom unitary [33]. [^5] We use “fundamental” to refer to the full quantum description of the gravitational system, as in [31]. In AdS/CFT, this assumption is valid whenever the time evolution of the CFT on a particular choice of code subspace is well approximated by a pseudorandom unitary at any finite value of $N$; [^6] We will eventually work at strictly infinite $N$, but then we will work only with time-evolved states rather than the unitary operator that generates that time evolution, since the aforementioned unitary operator is not well-defined in the strict infinite $N$ limit. this is not dissimilar from the assumption of e.g. [34] that the fundamental time evolution of a black hole is well modeled by a two-design or as in [31, 35] by a (pseudo)random unitary. Pseudorandom unitaries are operators drawn from an ensemble $\{U_{k}\}$ that are indistinguishable from Haar random unitaries by any efficient quantum algorithm. Our results below also hold for Haar random unitaries, but the advantage of pseudorandom unitaries over Haar random unitaries is that by definition they can be constructed by an efficient process. While existence of pseudorandom unitaries has not been explicitly demonstrated yet, candidate constructions have appeared in, e.g., [33]. We will not make use of any such explicit constructions.
@@
We now apply this result to our gravitational setup using the fact that reconstruction is a particular type of learning. We will consider code subspaces ${\cal H}_{\rm code}$ of dimension $e^{O(G_{N}^{a})}$, where $-1\leq a<0$, which do not necessarily have horizons in typical states. For example, a microcanonical window of width $O(N)$; another example is a set of states where the leading order contribution to the geometry is the same for all of the states at least for a time of $O(G_{N}^{0})$. We will show that it is easy to prepare candidates for such code subspaces without obviously incurring horizons. We next assume that the time evolution is gravitationally pseudorandom: we will define this term more precisely in Sec. 3.2, but roughly speaking it denotes an $O(G_{N}^{a})$ amount of pseudorandomness; we find that this amount of pseudorandomness is generally compatible with a ${\cal H}_{\rm code}$ of size $e^{O(G_{N}^{a})}$, which are the code subspaces that we consider. Next, we argue for the existence of an algorithm that, as part of the learning phase, can reconstruct the causal wedge of any efficiently-preparable state in this code subspace. Under certain technical assumptions, such an algorithm is able to output a guess $\widehat{U}$ for the time evolution of the initial state that accurately reproduces the expectation values of all operators in the causal wedge. However, the theorem of [32] guarantees that the fidelity between the predicted time evolution and the actual time evolved state is low, which implies that there exists a distinguishing operator $Q$ whose expectation values differ at $O(1)$ between the actual time evolved state and the predicted time evolution of the state.
@@
Let us now motivate the assumption of pseudorandomness as a model for singularity dynamics (which we have previously motivated for black holes; see also [35, 32]). A natural source of pseudorandom dynamics associated to a singularity is due to the Planckian physics at the singularity itself — what we described above as a UV-sensitive boundary condition. Let us briefly elaborate on this point. Intuitively, we may expect that to semiclassical observers — that is, to observers who are bounded in energy and complexity [31] — the radiation emitted from a timelike singularity or a spacelike singularity (with both past and future causal curves to the same $\mathscr{I}$, see Fig. 3(a)) appears roughly random. See Fig. 3(b) for a simple model.
Figure 4: (a) A spacelike singularity with both past and future causal curves to the same conformal boundary $\mathscr{I}$. (b) A simple model of pseudorandom evolution: radiation that is incoming at the naked singularity with some initial phase $\alpha_{\rm i}$ picks up a pseudorandom phase, such that the final phase $\alpha_{\rm f}$ is pseudorandomized.
Of course, the radiation cannot be actually Haar random (as this cannot be efficiently generated), so we instead expect that in many cases, pseudorandomness is an accurate model for the boundary dynamics describing a singularity. When this criterion is satisfied, the theorem of [32] together with the argument outlined above guarantees that the spacetime has an event horizon. [^12] Note that our argument does not guarantee that the singularity lies behind the event horizon, nor should it: we expect that small singularities can exist outside of horizons. Our techniques do suggest that it may be possible to limit the extent to which a singularity can lie outside of a horizon, but we leave a detailed investigation of this question to future work. It immediately follows that whenever the fundamental dynamics of a singularity are well approximated by a pseudorandom unitary, the singularity incurs an event horizon.
@@
We now briefly summarize the paper. We begin in the next subsection by setting up notations and conventions. In Sec. 2, we prove that nonminimal QESs must lie behind a horizon. In Sec. 3 we prove that when the time evolution in the fundamental description is sufficiently pseudorandom, the bulk spacetime contains an event horizon. We then apply this to the microcanonical window and a code subspace representing states with the same (leading order) geometry for times of $O(G_{N}^{0})$. In Sec. 4, we provide a definition that quantifies the size of (the naked part of) a singularity, and we show that classical (and semi-Planckian) singularities can accommodate gravitational pseudorandomness. By Cryptographic Censorship, this would then imply that such singularities must incur horizons. In Sec. 4, we also speculate on how Cryptographic Censorship can be used to prove a version of cosmic censorship. We end with a conjecture for a new form of cosmic censorship that is consistent with all known counterexamples to the standard WCCC. In Sec. 5, we end with a short discussion of our aspirations and a few possible future directions. In Appendices A and B, we provide proof details for the statements of Sec. 3.
@@ 1.1 Preliminaries
Unlike much of the AdS/CFT literature, we will not be imposing the assumption of global hyperbolicity on the bulk spacetimes under consideration, as this is tantamount to assuming weak (and strong) cosmic censorship. This means that, even in the $G_{N}\rightarrow 0$ limit, QFT in curved spacetime (including linearized gravitons) cannot be utilized for time evolution. However, we will still assume that QFT in curved spacetime is valid within causal diamonds that contain no singularities. We will further assume that any spacetime under consideration admits an initial data slice $\Sigma$ with an everywhere defined and nowhere vanishing continuous timelike unit normal; in particular, we require that $D[\Sigma]$ contains an open set containing $\Sigma$. Any singularity in the spacetime will be assumed to act as a UV-sensitive boundary condition that determines the relation between the algebras of different causal diamonds. [^13] The singularity only has an effect on the relation between two diamonds $D_{1}$ and $D_{2}$ when the set $[J^{+}(D_{1})\cup J^{-}(D_{1})]\cap[J^{+}(D_{2})\cup J^{-}(D_{2})]$ contains part of the singularity. If this set does not contain a part of the singularity, the EFT can be used to relate the algebra of observables in $D_{1}$ and $D_{2}$. Finally, we will take a conformal frame which is taken to be fixed at any value of $N$, including in the large-$N$ limit.
@@ 3.2 Complexity of learning in holography
    for any $\widehat{U}\leftarrow\mathcal{A}^{U}$, there exists an operator $Q$ (which may depend on $U$, |$\ket{\psi}$, and $\widehat{U}$) such that
@@ 3.3 Pseudorandomness implies horizons
Point (b) is in a sense obvious holographically given the developments on complexity of reconstruction in AdS/CFT. The general expectation following from [30] is that a simple quantum algorithm should be able to reconstruct the causal wedge. However, as our presentation thus far has been at a higher level of rigor, we aim to illustrate more precisely that under a mild technical assumption, there exists a quantum learning algorithm satisfying the assumptions of Theorem 2 that, given oracle access to $U$, can output an operator $V$ where $V\ket{\psi}$ can accurately reproduce the expectation values of all simple operators in the state $U\ket{\psi}$. Applicability to our strengthened version, Theorem 3, does not immediately follow and would require techniques from measure concentration. This falls outside the scope of the current work, but we expect that this may go through if the algorithm in [61] satisfies certain assumptions. Readers willing to accept the existence of the relevant algorithm — i.e., an efficient algorithm for reconstructing the causal wedge — can skip directly to the statement of Cryptographic Censorship, Theorem 5 below.
@@ 4.1 The size of singularities
If |$\ket{\psi(t)}$ is the state whose complete past and future history is dual to $(M,g)$, then without access to the time evolution operator, we can only reconstruct operators at points in spacetime that are not causally related to the boundary time slice on which |$\ket{\psi(t)}$ lives — i.e., the Wheeler–de Witt patch $W_{t}$. [^37] This assumes for simplicity that there is only one asymptotic boundary, and that the state of the CFT on that boundary is pure — so the entanglement wedge of |$\ket{\psi}$ is the entire bulk. If there are multiple boundaries or the state on one boundary is mixed for a different reason, we simply take the intersection of $W_{t}$ with the entanglement wedge of |$\ket{\psi}$. Let us now specialize to a spacetime with a naked singularity $\Gamma$. For some fixed $t$, generally only a (possibly empty) subset of $\Gamma$ will be contained in $W_{t}$: to learn everything about $\Gamma$ (for example by reconstruction of all near-singularity correlators), a certain amount of boundary time evolution is necessary. More precisely, if $T_{-}$ is nonempty, and we are evolving forwards from the past, then the past-most boundary of $T_{-}$ corresponds to the earliest time that some part of $\Gamma$ is encoded in the Wheeler–de Witt patch $W_{t}$. As we evolve forward through $T_{-}$, new pieces of $\Gamma$ enter $W_{t}$ and are encoded in the time-evolved state. Finally, once we evolve past the futuremost boundary of $T_{-}$, no new parts of $\Gamma$ enter $W_{t}$ that we have not already seen, and so further forward time evolution reveals nothing new about the singularity. Thus, everything about the naked singularity can be learned by carrying out CFT time evolution through the band $T_{-}$ (plus bulk reconstruction not using the time-evolution operator). For a singularity where $T_{+}$ is also nonempty, everything about the singularity could alternatively have been learned by evolving through $T_{+}$. So as advertised, $|\Gamma|_{\mathscr{I}}=\min\{|T_{-}|,|T_{+}|\}$ measures the minimal amount of CFT time evolution required to learn everything about the naked singularity. Note that $|\Gamma|_{\mathscr{I}}$ has units of spacetime volume rather than time. This is a feature: if all information about the singularity is available upon evolution of the state forwards in a small spatial subregion of a boundary Cauchy slice, then this should be accounted for in the smallness of a singularity. See Fig. 11. Note also that if $|\Gamma|_{\mathscr{I}}$ is large, this always is caused by a large amount of time evolution, since timeslices have bounded volume.

### 19 Mar 2024 v2

@@ 1 Introduction
Curvature singularities are ubiquitous in General Relativity: the evolution of smooth initial data generically results in such singularities [1, 2, 3]. Observationally, however, evidence for singularity formation from generic gravitational processes (such as matter collapse) remains at best indirect. This apparent absence of directly-observable singularities is often explained by the general expectation that the formation of horizons is likewise generic: so singularities are usually cloaked behind event horizons. This hypothesis — that singularities resulting from initial data evolution should be generically hidden behind horizons — is known as the Weak Cosmic Censorship Conjecture (WCCC) [4]. It is a cornerstone assumption of almost all results in modern gravitational physics, from classical theorems such as the Hawking area law [5], to key results from the general “it from qubit” program. In particular, the expectation that CFT thermalization corresponds to black hole equilibration, and more generally that black hole statistics are reflected in CFT thermal states, crucially depends on the assumption that states with horizons are typical in some measure. [^1] This is not identical to the firewall sense of typicality of [6]. See footnote 6 on page 6. If the WCCC were false, naked singularities could be as typical as black holes; our understanding of quantum chaos, eigenstate thermalization, and black hole information in holography would be called into question.
@@ A General Guarantee for Horizon Formation
More concretely, we ask when there is a guarantee for the existence of some operator $Q$ in the large-$N$ limit that cannot be reconstructed from knowledge of the causal wedge alone, given that the latter can be reconstructed efficiently, e.g. by jumping into the bulk, making a measurement, and sailing back towards the boundary. [^4] By an efficient reconstruction, we mean one that can be implemented by a quantum circuit whose size is polynomially bounded in the logarithm of the dimension of the code subspace under consideration.
In [32], two of us provided such a guarantee at large but finite $N$ under certain assumptions. The full statement of the relevant theorem will be reviewed at length in Sec. 3.1; here we provide a heuristic statement only, starting with the most significant assumption: that the action of the fundamental time evolution operator $U$ of the system is well modeled on our code subspace by a pseudorandom unitary [33] (we use “fundamental” to refer to the full quantum description of the gravitational system, as in [31]). In AdS/CFT, this assumption is valid whenever the time evolution of the CFT on a particular choice of code subspace is well approximated by a pseudorandom unitary at any finite value of $N$; [^5] We will eventually work at strictly infinite $N$, but then we will work only with time-evolved states rather than the unitary operator that generates that time evolution, since the aforementioned unitary operator is not well-defined in the strict infinite $N$ limit. this is not dissimilar from the assumption of e.g. [34] that the fundamental time evolution of a black hole is well modeled by a two-design or as in [31, 35] by a (pseudo)random unitary. Pseudorandom unitaries are operators drawn from an ensemble $\{U_{k}\}$ that are indistinguishable from Haar random unitaries by any efficient quantum algorithm. Our results below also hold for Haar random unitaries, but the advantage of pseudorandom unitaries over Haar random unitaries is that by definition they can be constructed by an efficient process. While existence of pseudorandom unitaries has not been explicitly demonstrated yet, candidate constructions have appeared in, e.g., [33]. We will not make use of any such explicit constructions.
@@
We now apply this result to our gravitational setup using the fact that reconstruction is a particular type of learning. We will consider code subspaces ${\cal H}_{\rm code}$ of dimension $e^{O(G_{N}^{a})}$, where $-1\leq a<0$, which do not necessarily have horizons in typical states. For example, a microcanonical window of width $O(N)$; another example is a set of states where the leading order contribution to the geometry is the same for all of the states at least for a time of $O(G_{N}^{0})$. We will show that it is easy to prepare candidates for such code subspaces without obviously incurring horizons. We next assume that the time evolution is gravitationally pseudorandom: we will define this term more precisely in Sec. 3.2, but roughly speaking it denotes an $O(G_{N}^{a})$ amount of pseudorandomness; we find that this amount of pseudorandomness is generally compatible with an ${\cal H}_{\rm code}$ of size $e^{O(G_{N}^{a})}$, which are the code subspaces that we consider. Next, we argue for the existence of an efficient algorithm that, as part of the learning phase, can reconstruct the causal wedge of any efficiently-preparable state in this code subspace. Under certain technical assumptions, there is an algorithm that, given access to some CFT data, is able to output a guess $\widehat{U}$ for the time evolution of the initial state that accurately reproduces the expectation values of all operators in the causal wedge. However, the theorem of [32] guarantees that the fidelity between the predicted time evolution and the actual time evolved state is low, which implies that there exists a distinguishing operator $Q$ whose expectation values differ at $O(1)$ between the actual time evolved state and the predicted time evolution of the state.
@@
Let us now motivate the assumption of pseudorandomness as a model for singularity dynamics (which we have previously motivated for black holes; see also [35, 32]). A natural source of pseudorandom dynamics associated to a singularity is due to the Planckian physics at the singularity itself — what we described above as a UV-sensitive boundary condition. We expect such a boundary condition cannot be predicted by effective field theory. We will argue in more detail in Sec. 4.2 that this UV-sensitive boundary condition at the singularity could plausibly be the effect of pseudorandomness in the fundamental dynamics. We expect that in many cases, pseudorandomness is an accurate model for the boundary dynamics describing a singularity. When this criterion is satisfied, the theorem of [32] together with the argument outlined above guarantees that the spacetime has an event horizon. [^11] Note that our argument does not guarantee that the singularity lies behind the event horizon, nor should it: we expect that small singularities can exist outside of horizons. Our techniques do suggest that it may be possible to limit the extent to which a singularity can lie outside of a horizon, but we leave a detailed investigation of this question to future work. It immediately follows that whenever the fundamental dynamics of a singularity are well approximated by a pseudorandom unitary, the singularity incurs an event horizon.
@@
We emphasize here that throughout this paper we will be considering pseudorandomness in the boundary dynamics. Readers interested in understanding what happens in the situation of bulk pseudorandomness should refer to Sec. 5.
We now briefly summarize the paper. We begin in the next subsection by setting up notations and conventions. In Sec. 2, we prove that nonminimal QESs must lie behind a horizon. In Sec. 3 we prove that when the time evolution in the fundamental description is sufficiently pseudorandom, the bulk spacetime contains an event horizon. We then apply this to the microcanonical window and a code subspace representing states with the same (leading order) geometry for times of $O(G_{N}^{0})$. In Sec. 4, we provide a definition that quantifies the size of (the naked part of) a singularity, and we show that classical (and semi-Planckian) singularities can accommodate gravitational pseudorandomness. By Cryptographic Censorship, this would then imply that such singularities must incur horizons. In Sec. 4, we also speculate on how Cryptographic Censorship can be used to prove a version of cosmic censorship. We end with a conjecture for a new, inherently quantum gravitational form of cosmic censorship that may be consistent with known counterexamples to the standard WCCC, including the evaporating black hole. In Sec. 5, we end with a short discussion of our aspirations and a few possible future directions. In Apps. A and B, we provide proof details for the statements of Sec. 3.
@@ 1.1 Preliminaries
Unlike much of the AdS/CFT literature, we will not be imposing the assumption of global hyperbolicity on the bulk spacetimes under consideration, as this is tantamount to assuming weak (and strong) cosmic censorship. This means that, even in the $G_{N}\rightarrow 0$ limit, QFT in curved spacetime (including linearized gravitons) cannot be utilized for time evolution. However, we will still assume that QFT in curved spacetime is valid within causal diamonds that contain no singularities. We will work with a code subspace of states in the CFT and assume that the algebra of CFT operators at strictly infinite $N$ can be identified with the algebra of QFT operators on the spacetime emergent at infinite $N$ (see [49] for a discussion of these algebras). We will further assume that any spacetime under consideration admits an initial data slice $\Sigma$ with an everywhere defined and nowhere vanishing continuous timelike unit normal; in particular, we require that $D[\Sigma]$ contains an open set containing $\Sigma$. Any singularity in the spacetime will be assumed to act as a UV-sensitive boundary condition that determines the relation between the algebras of different causal diamonds. [^12] The singularity only has an effect on the relation between two diamonds $D_{1}$ and $D_{2}$ when the set $[J^{+}(D_{1})\cup J^{-}(D_{1})]\cap[J^{+}(D_{2})\cup J^{-}(D_{2})]$ contains part of the singularity. If this set does not contain a part of the singularity, the EFT can be used to relate the algebra of observables in $D_{1}$ and $D_{2}$. Finally, we will take a conformal frame which is taken to be fixed at any value of $N$, including in the large-$N$ limit.
@@ 3.2 Complexity of learning in holography
    there exists an operator $Q$ (which may depend on $U$, |$\ket{\psi}$, and $\widehat{U}$) such that
@@ 3.3 Pseudorandomness implies horizons
Point (b) is in a sense obvious holographically given the developments on complexity of reconstruction in AdS/CFT. The general expectation following from [30] is that a simple quantum algorithm should be able to reconstruct the causal wedge. We might worry that the existence of a naked singularity complicates this story, but there is a simple physical argument suggesting otherwise. To measure an operator in the causal wedge, we can send rocket ships from the conformal boundary that travel into the bulk, carry out measurements anywhere in the causal wedge, and then travel back out to the boundary with the result. Since we are only interested in measuring operators that survive at strictly infinite $N$, the complexity to implement these rocket ships from the CFT should not scale exponentially with $N$ [^26] One might worry that the bulk measurements take too long for the rocket ships to be able to get back out to the boundary. Again, this should not be the case for bulk operators that survive the infinite $N$ limit. Under the assumption that there exists an efficient process that can produce a guess $V$ for the time evolution operator $U$ using these measurements, we can then combine this with point (a) above to formulate our theorem of Cryptographic Censorship, Theorem 5 below.
Figure 7: In the leftmost and middle panels, the reconstruction of operators inside the green wedge (which is globally hyperbolic) can be done using the HKLL protocol, which should be simple, insofar as we can compute the expectation values of local operators in the boundary theory on the time band $\Delta t$. We can keep doing this for subsequent time bands until we have reconstructed the full causal wedge, as in the rightmost panel. We use this as an ingredient to motivate the existence of an algorithm that can output an operator $V$ which reproduces the time evolution of operators in the causal wedge.
In many cases, we can motivate this assumption — the existence of an efficient learning algorithm that can produce such a guess $V$ — from a computer science point of view. For instance, consider first a domain of dependence corresponding to a boundary time band. If we have access to the time-evolved expectation values of single-trace operators in that time band, we can use the HKLL protocol [56, 57, 58] to reconstruct this domain of dependence. See Fig. 7. In the case where the causal wedge can be generated from a union of such wedges, and if there exists a simple algorithm that can aggregate all these reconstructed wedges into a guess for the time evolution operator, then we can efficiently reconstruct the full causal wedge. There is in fact a learning algorithm that comes close to providing such a guess $V$, as we will explain in more detail below; this thus provides strong motivation that learning the causal wedge is simple, even in the presence of such naked singularities.
However, it is possible that there exist naked singularities with unusual topologies where this reasoning falls short and we cannot use this ‘progressive HKLL’ protocol combined with the aforementioned existing learning algorithm. An illustration of such a hypothetical singularity is given in Fig. 8, where the interior of a cylindrical singularity does not lie inside any HKLL wedge. In such cases, however, we revert to our initial holographic reasoning to motivate our assumption: we can still use rocket ships to measure operators in the entire causal wedge.
For the case where HKLL in progressive wedges does work, we will now illustrate more precisely that under a mild technical assumption, there exists a quantum learning algorithm satisfying the assumptions of Theorem 2 that, given oracle access to $U$, can output an operator $V$ where $V\ket{\psi}$ can accurately reproduce the expectation values of all simple operators in the state $U\ket{\psi}$. Applicability to our strengthened version, Theorem 3, does not immediately follow and would require techniques from measure concentration. This falls outside the scope of the current work, but we expect that this may go through if the algorithm in [62] satisfies certain assumptions. Readers willing to accept the existence of the relevant algorithm — i.e., an efficient algorithm for reconstructing the causal wedge — can skip directly to the statement of Cryptographic Censorship, Theorem 5 below.
Figure 8: The region of spacetime inside this cylindrical naked singularity does not lie inside of any HKLL wedge. It is possible, however, for a rocket ship to enter the region behind the singularity and come back out to the boundary.
@@ 4.1 The size of singularities
If |$\ket{\psi(t)}$ is the state whose complete past and future history is dual to $(M,g)$, then without access to the time evolution operator, we can only reconstruct operators at points in spacetime that are not causally related to the boundary time slice on which |$\ket{\psi(t)}$ lives — i.e., the Wheeler–de Witt patch $W_{t}$. [^37] This assumes for simplicity that there is only one asymptotic boundary, and that the state of the CFT on that boundary is pure — so the entanglement wedge of |$\ket{\psi}$ is the entire bulk. If there are multiple boundaries or the state on one boundary is mixed for a different reason, we simply take the intersection of $W_{t}$ with the entanglement wedge of |$\ket{\psi}$. Let us now specialize to a spacetime with a naked singularity $\Gamma$. For some fixed $t$, generally only a (possibly empty) subset of $\Gamma$ will be contained in $W_{t}$: to learn everything about $\Gamma$ (for example by reconstruction of all near-singularity correlators), a certain amount of boundary time evolution is necessary. More precisely, if $T_{-}$ is nonempty, and we are evolving forwards from the past, then the past-most boundary of $T_{-}$ corresponds to the earliest time that some part of $\Gamma$ is encoded in the Wheeler–de Witt patch $W_{t}$. As we evolve forward through $T_{-}$, new pieces of $\Gamma$ enter $W_{t}$ and are encoded in the time-evolved state. Finally, once we evolve past the futuremost boundary of $T_{-}$, no new parts of $\Gamma$ enter $W_{t}$ that we have not already seen, and so further forward time evolution reveals nothing new about the singularity.
@@
Thus, everything about the naked singularity can be learned by carrying out CFT time evolution through the band $T_{-}$ (plus bulk reconstruction not using the time-evolution operator). For a singularity where $T_{+}$ is also nonempty, everything about the singularity could alternatively have been learned by evolving through $T_{+}$. So as advertised, $|\Gamma|_{\mathscr{I}}=\min\{|T_{-}|,|T_{+}|\}$ measures the minimal amount of CFT time evolution required to learn everything about the naked singularity. Note that $|\Gamma|_{\mathscr{I}}$ has units of spacetime volume rather than time. This is a feature: if all information about the singularity is available upon evolution of the state forwards in a small spatial subregion of a boundary Cauchy slice, then this should be accounted for in the smallness of a singularity. See Fig. 12. Note also that if $|\Gamma|_{\mathscr{I}}$ is large, this always is caused by a large amount of time evolution, since timeslices have bounded volume.